# Diffusion-Driven Semantic Communication for Generative Models with Bandwidth Constraints

### v1

@@ V-C Performance of Compression with Different Bandwidths
In this subsection, we evaluate the performance of the proposed generation semantic communication with compression and guidance across various compression rates. The JSCC system is used as our benchmark for performance comparison. Since previous studies have demonstrated that DeepJSCC outperforms classical separation-based methods, we exclude them from consideration as baselines.
In the figures, the legends with “Proposed ($C$)” and “DJSCC ($C$)” represent our proposed generation semantic communication system and the DJSCC system with different compression rates, respectively. In the DJSCC scheme, the same encoder and decoder network structures as ours are employed, and guidance from the uncompressed network is also incorporated. However, reparameterization is not utilized to ensure a Gaussian distribution in the output. Additionally, since the DJSCC scheme is substituted in the compressed generator, the loss function is also modified to the mean squared error (MSE) between the outputs of the uncompressed generator and the compressed generator. Here, $C$ denotes the channel dimensionality of the transmitted features, where fewer channels indicate higher compression rates. The mapping between compression rate $r$ and $C$ is defined as $r=0.0013C$. Our objective is to ensure accurate transmission of semantic information while using less bandwidth.
Fig. 6 shows the comparisons in terms of PSNR, SSIM, LPIPS and CLIP-score metrics, respectively. The results show that our proposed generation semantic communication system has significant performance improvement compared to the DJSCC at all evaluated values of SNR and the compression rate.
V-D Improvement from Reparameterization and Guidance
Guidance provides auxiliary information derived from the uncompressed generator, which guides the upsampling network to learn the distribution of the uncompressed generator, directing the output of the upsampling model towards more accurate distributions. In this subsection, we analyze the effects of guidance on enhancing the performance of the generation semantic communication based on diffusion. The results in Fig. 7 show improvements for two compression rates, namely $C=14$ and $C=4$. Compared to DJSCC, the generation semantic communication system demonstrates improvements in two aspects. One part of the improvement stems from maintaining the Gaussian distribution of features through reparameterization, while the other part arises from the guidance provided by the uncompressed generator. In terms of pixel-level metrics such as PSNR and SSIM, the performance improvement mainly comes from reparameterization, while for semantic metrics like LPIPS and CLIP-score, the enhancement primarily comes from guidance. Therefore, overall, both reparameterization and guidance from the uncompressed generator play crucial roles in enhancing transmission performance.
@@ V-E Hyperparameters Settings in the Hybrid Loss
TABLE V: The performance of different $\lambda$ in LSUN-Churches datasets, when $\gamma=0$ and SNR=5dB.
@@
TABLE VI: The performance of different $\gamma$ in LSUN-Churches datasets, when $\lambda=0.1$ and SNR=5dB.
@@
Here, our focus lies on configuring the hyperparameters within the loss function for compression and guidance, which is employed to train the diffusion model. Hyperparameters, predetermined variables set before the learning process begins and remain constant during training, play a pivotal role in balancing the diverse components of the loss function and can significantly impact the capacity of model to learn from data. We investigate various hyperparameter configurations and their implications on performance. To conserve training resources, we subset the training dataset to include only 1000 images and evaluate it on a testing dataset consisting of 300 images. Training for 20 epochs allows us to identify the optimal combination based on selected metrics. Initially, we set the hyperparameter $\gamma$ to 0 to determine the optimal value of the hyperparameter $\lambda$. Subsequently, after identifying the best value for $\lambda$, we adjust $\gamma$ to select the optimal combination. As depicted in Tab. V and Tab. VI, we ultimately selected the parameter combination of $\gamma=0.1$ and $\lambda=0.1$, which exhibited the best performance.
@@ VI Conclusion
In this paper, we introduced a novel diffusion-driven semantic communication framework that incorporates advanced VAE-based compression for generative model in bandwidth-constrained environments. By leveraging the diffusion model, where the signal transmission process over the wireless channel serves as the forward diffusion process, our architecture successfully eliminates Gaussian noise from channel while conserving bandwidth, enhancing its applicability in wireless communication scenarios. We integrated a downsampling module and a corresponding VAE-based upsampling module with reparameterization to ensure the recovered features conform to the Gaussian distribution, a prerequisite for the diffusion model. Furthermore, we derived the loss function with the guidance for this system based on the design of reparameterization and compression. Experimental results demonstrate the effectiveness of our approach, with improvements observed in both compression rates and SNR compared to the DJSCC baseline. Specifically, the reparameterization significantly enhances pixel-level performance, while guidance plays a crucial role in improving semantic transmission. This comprehensive approach demonstrates the potential for advancing semantic communication systems, offering enhanced reliability in generation performance.

### v2

@@ V-C Performance of Compression with Different Channels and Bandwidths
In this subsection, we evaluate the performance of the proposed generative semantic communication framework with compression and guidance across various compression rates. The method “CDDM” [20], which integrates diffusion models with the DeepJSCC network, is employed as our benchmark for performance comparison. While CDDM uses the same encoder and decoder network structures and incorporates guidance from the pre-trained diffusion model with the MSE loss function, our approach differs by introducing a VAE-based compression module for bandwidth adaptation and a hybrid loss function to further enhance reconstruction quality. Since previous studies have shown that DeepJSCC consistently outperforms classical separation-based methods, we exclude those methods from consideration as baselines.
TABLE V: Experimental requirements of an epoch on the LSUN-Churches dataset.
Training time (Memory usage) 34 h (27139 MiB)
Inference time (Memory usage) 7 s (26125 MiB)
We compare the performance of our method and CDDM under different channels. Additionally, we further evaluate and compare their performance under AWGN channels across various bandwidth constraints. Fig. 6 and Fig. 7 present the comparisons in terms of PSNR, SSIM, LPIPS, and CLIP-score metrics, respectively. As shown in Fig. 6, our proposed method achieves better performance across all evaluated channels. In the Fig. 7, the legends with “Proposed ($C$)” and “CDDM ($C$)” represent our proposed generation semantic communication system and the CDDM system with different compression rates, respectively. Here, $C$ denotes the channel dimensionality of the transmitted features, where fewer channels indicate higher compression rates. The mapping between compression rate $r$ and $C$ is defined as $r=0.0013C$. Our objective is to ensure accurate transmission of semantic information while using less bandwidth. Fig. 7 shows the comparisons in terms of PSNR, SSIM, LPIPS and CLIP-score metrics, respectively. The results show that our proposed generation semantic communication system has significant performance improvement compared to the CDDM at all evaluated values of SNR and the compression rate. Further, Table V presents the training and inference times, as well as the memory requirements of the proposed methods on the LSUN-Churches dataset. Specifically, the inference time is evaluated under 10 dB SNR, the reverse steps equivalently converted to 70.
V-D Ablation study
TABLE VI: The ablation of the reparameterization and the proposed loss on LSUN-Churches dataset across different SNRs.
Metric Methods SNR
VAE $L_{\text{MSE}}$ $L_{\text{KL}}$ $L_{g}$ 1dB 3dB 5dB 7dB 9dB 12dB
PSNR $\uparrow$ 13.79 14.39 14.56 14.82 14.96 15.02
16.15 17.51 18.32 18.72 18.94 19.09
15.78 17.06 17.73 17.92 17.95 18.02
16.66 18.25 19.28 19.91 20.35 20.64
16.26 17.99 19.21 20.01 20.50 20.71
16.61 18.31 19.45 20.15 20.55 20.76
SSIM $\uparrow$ 0.35 0.35 0.35 0.36 0.37 0.37
0.43 0.45 0.47 0.48 0.49 0.49
0.40 0.43 0.44 0.44 0.45 0.45
0.45 0.48 0.51 0.53 0.55 0.56
0.45 0.49 0.52 0.54 0.56 0.56
0.46 0.49 0.52 0.54 0.56 0.57
CLIP $\uparrow$ 0.45 0.48 0.49 0.50 0.53 0.54
0.50 0.51 0.52 0.53 0.56 0.58
0.55 0.58 0.59 0.62 0.66 0.70
0.54 0.56 0.58 0.61 0.64 0.69
0.56 0.58 0.60 0.64 0.68 0.72
0.56 0.58 0.61 0.65 0.70 0.74
FID $\downarrow$ 268.89 254.21 244.21 217.34 206.88 193.05
249.67 238.40 230.23 217.34 197.17 168.68
207.39 182.78 156.71 122.33 91.50 71.78
210.74 191.18 169.47 135.49 102.14 77.67
196.70 172.22 142.69 108.16 78.06 61.77
193.54 168.72 139.46 102.83 71.39 59.87
LPIPS $\downarrow$ 0.70 0.69 0.69 0.68 0.68 0.68
0.69 0.69 0.69 0.68 0.68 0.67
0.64 0.60 0.57 0.53 0.49 0.47
0.64 0.62 0.60 0.56 0.53 0.50
0.63 0.60 0.56 0.51 0.48 0.45
0.57 0.53 0.47 0.45 0.43 0.42
To comprehensively evaluate the contributions of various components in our proposed framework, ablation experiment is provided under different SNR conditions. The results are summarized in Table Table VI, where each row represents the performance of the system using a different combination of components, including the VAE module and specific loss terms, such as $L_{\text{MSE}}$ and $L_{\text{KL}}$ of $L_{v}$ and $L_{g}$. Metrics such as PSNR, SSIM, CLIP-score, FID, and LPIPS are employed to evaluate both the pixel-level reconstruction fidelity and the semantic-level consistency of the generated results.
The baseline method, which employs only $L_{\text{MSE}}$ without the VAE module, demonstrates the weakest performance across all metrics, particularly under low SNR conditions. Incorporating the VAE module to maintain Gaussian-distributed features results in significant improvements in PSNR and SSIM, reflecting enhanced fidelity in the reconstructed images. The inclusion of $L_{\text{KL}}$ further boosts performance, as it contributes to better regularization and consistency in the latent space representation. Additionally, the integration of $L_{g}$ , which aligns the semantic content with the guidance provided by the pre-trained diffusion model, consistently enhances all evaluated metrics. Overall, the integration of all components ($L_{\text{MSE}}$, $L_{\text{KL}}$, $L_{g}$, and the VAE module) achieves the best performance, as evidenced by the highest PSNR (20.76), SSIM (0.57), and CLIP score (0.74), alongside the lowest FID (59.87) and LPIPS (0.42). These results demonstrate the effectiveness of the proposed framework in both semantic-level and pixel-level evaluations. Each component contributes to the overall system performance, and their combined integration leads to an improvement in both semantic preservation and detailed reconstruction quality under bandwidth constraints.
@@ V-E Hyperparameters Settings in the Hybrid Loss
TABLE VII: The performance of different $\lambda$ in LSUN-Churches datasets, when $\gamma=0$ and SNR=5dB.
@@
TABLE VIII: The performance of different $\gamma$ in LSUN-Churches datasets, when $\lambda=0.1$ and SNR=5dB.
@@
Here, our focus lies on configuring the hyperparameters within the loss function for compression and guidance, which is employed to train the diffusion model. Hyperparameters, predetermined variables set before the learning process begins and remain constant during training, play a pivotal role in balancing the diverse components of the loss function and can significantly impact the capacity of model to learn from data. We investigate various hyperparameter configurations and their implications on performance. To conserve training resources, we subset the training dataset to include only 1000 images and evaluate it on a testing dataset consisting of 300 images. Training for 20 epochs allows us to identify the optimal combination based on selected metrics. Initially, we set the hyperparameter $\gamma$ to 0 to determine the optimal value of the hyperparameter $\lambda$. Subsequently, after identifying the best value for $\lambda$, we adjust $\gamma$ to select the optimal combination. As depicted in Tab. VII and Tab. VIII, we ultimately selected the parameter combination of $\gamma=0.1$ and $\lambda=0.1$, which exhibited the best performance.
@@ VI Conclusion
In this paper, we introduced a novel diffusion-driven semantic communication framework that incorporates advanced VAE-based compression for generative model in bandwidth-constrained environments. By leveraging the diffusion model, where the signal transmission process over the wireless channel serves as the forward diffusion process, our architecture successfully eliminates Gaussian noise from channel while conserving bandwidth, enhancing its applicability in wireless communication scenarios. We integrated a downsampling module and a corresponding VAE-based upsampling module with reparameterization to ensure the recovered features conform to the Gaussian distribution, a prerequisite for the diffusion model. Furthermore, we derived the loss function with the guidance for this system based on the design of reparameterization and compression. Experimental results demonstrate the effectiveness of our approach, with improvements observed in both compression rates and SNR compared to the CDDM baseline. Specifically, the reparameterization significantly enhances pixel-level performance, while guidance plays a crucial role in improving semantic transmission. This comprehensive approach demonstrates the potential for advancing semantic communication systems, offering enhanced reliability in generation performance.